• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Meziani

    Log entry time 00:05:01 on June10,2004

    Entry number 125203

    keyword=Shift Summary

    Swing shift started with Hall A in controlled access

    16:00 Hall A in controlled access.
    18:50 Target in empty position should not be moved (D. Meekins)
    Valves opened and vaccum about 610^-6 in target chamber and spectrometer

    19:30 D. L'huillier finished working on the diode system on the left arm.

    20:30 Beam back. Oleksander Glamazdin is here for a Moller measurement
    23:00 Moeller measurement done.
    MCC is tuning beam through the Compton chicane
    requested positions IPM1P02A x = -0.50um y=-0.85
    IPM1P02B x = -0.50um y=-0.85
    24:00 MCC still tuning beam